# Stability of Jordan Recurrent Neural Network Estimator

### v1

@@ Abstract
State estimation involves finding the states of a noisy dynamical system with a noisy initial condition using noisy measurements, given a model of the system. Many methods exist for state estimation of dynamical systems. For linear dynamical systems with white Gaussian noises of known mean and variance, Kalman filtering is an optimal state estimation algorithm. While several extensions of the Kalman filter to nonlinear systems exist, either the method is not optimal or calculation is not feasible. Several approaches using neural networks have been developed in recent years. While these approaches seem to work numerically, stability and convergence guarantees do not exist. In this paper, we propose to use a Jordan recurrent neural network (JRN) for state estimation. An input-to-state stability (ISS) analysis of the error dynamics is used. Our results are compared to the Kalman filter for the linear case and the extended Kalman filter for the nonlinear case.
@@ I INTRODUCTION
State estimation refers to estimating the states of a noisy dynamical system given noisy and incomplete measurements, partial information about the initial state, as well as a mathematical model for the system. In practical applications, estimating the state of a system is a challenging task for various reasons. The complexity of high-order systems, the availability of a finite number of measurements, inaccurate measurements, incomplete information about the initial state, and errors in the model can make it difficult to estimate the state of the system.
One of the most commonly used optimal state estimation methods is the Kalman filter (KF) [1]. It is optimal for linear systems with white Gaussian noises with zero mean and known covariance. When implemented for a discrete-time system, it consists of two steps: prediction and filtering, and is a recursive algorithm. Many researchers have extended the Kalman filter to nonlinear systems. The most obvious model-based generalization of this filter to nonlinear systems is the Extended Kalman filter (EKF) [2] [3]. The EKF relies on linearization of the dynamical system at each time-step. Although local convergence guarantees [4, 5] exist, convergence to the true state is an issue for highly nonlinear systems and when the estimate of the initial condition is very inaccurate. Moreover, even when the estimation error converges to zero, the method is not optimal.
In recent years, machine learning approaches, particularly artificial neural networks, have attracted attention. Gao et al. [6] used deep long short-term memory (DLSTM) networks to approximate the predicted and filtered state from observations. Although they aimed to study trajectory tracking, their approach turned out to be a state estimation problem since they ignored certain assumptions and used knowledge of the model for training dataset generation. They presented two frameworks, Bayesian and non-Bayesian. For the Bayesian approach, they used Bayes’ rule [7] and designed the DLSTM so that it predicts and filters iteratively within the same network. For the non-Bayesian approach, they considered two DLSTM networks, one for prediction and the other for filtering, based on the fact that neural networks can perform both steps independently. Thus, one network performed prediction without measurements, and the other corrected the predictions by using new measurements. Estimation accuracy, error bounds, analytic performance evaluation, and neural network convergence remain open problems for these approaches. Breitan et al. [8] proposed a machine learning framework that overcame the requirement to find solutions to the Hamilton-Jacobi-Bellman (HJB) partial differential equation for an optimal estimator in the continuous-time case. A neural network was used to approximate the gradient of the solution of the HJB equation. They exploited the HJB equation to construct the observer gain, which coincided with the Kalman-Bucy filter for a linear system. Peralez [9] further proposed a deep learning-based Luenberger estimator design for discrete-time nonlinear systems by approximating the mapping using unsupervised learning. Gençay et al [10] compared the method of least squares, Elman recurrent neural network (ERN), and feedforward neural networks (FNN) for noisy time series data with fully observable states. They concluded that ERNs provided the best estimation properties among the three.
Figure 1: The structure of the Jordan recurrent network proposed for state estimation. It uses output-to-hidden recurrent connections similar to the filter’s dynamical system. The symbols $y^{(t)},a^{(t)},x^{(t)}$ and $\hat{x}^{(t)}$ represent input measurement vector, hidden layer vector, true state vector and estimated state vector respectively at time $t$. Cost function $J^{(t)}$ is the mean squared error at time $t$. The weights and biases of the network are represented by $W_{ay},W_{ax},W_{xa},b_{x}$ and $b_{y}.$
Identifying states entirely on the basis of input-output data is called system identification. Recurrent neural networks have been shown to work well for system identification [11]. In fact, Jordan recurrent networks (JRNs) have been used for system identification by several researchers and have been shown to work as well as ERNs [12][13]. A convergence analysis, with some assumptions, of both networks for system identification has been discussed in [14]. Forecasting is a domain in which future values are predicted based on observed trends. In [15], an Elman-based long short-term memory (LSTM) network is compared with an EKF for time series forecasting. The authors conclude that the LSTM performs better than the EKF for nonlinear systems. In [16], authors show that the ERN performs at least as well as FNN for Lyapunov exponents forecasting, discussing cases where it performs better in detail as well.
This encouraged us to use recurrent neural networks for state estimation. The issue with ERNs is that proving the stability of the estimator is difficult, as the dependence is not directly on the estimated state but on the hidden state at the previous time-step. This is not the case for a JRN. We use an algorithm that uses known information about a dynamical system to train a JRN for state estimation. The structure of JRNs resembles that of a dynamical system, making it a suitable candidate for state estimation. The difference between Elman and Jordan recurrent neural networks lies in the types of recurrent connections [17]. Schäfer and Zimmerman [18] proved ERNs are universal approximators of dynamical systems. These results can be extended to JRNs for state estimation [19].
@@
The stability of recurrent neural networks (RNNs) in the sense of Lyapunov has been well-developed in [20, 21], although the stability of errors in RNNs as neural estimators for state estimation has not been extensively explored. Finding Lyapunov functions is generally challenging. Several papers [22, 23] have shown that Lyapunov functions can be attained and represented by expressions of the compositional structure of the neural network, the correctness being guaranteed by satisfiability modulo theories (SMT) solvers. In this paper, we design an estimator using JRNs and employ an input-to-state stability (ISS) approach to analyze the stability of the error dynamics. We provide a theoretical result establishing a condition under which the error dynamics will converge to zero if the original system is stable. Estimation of unstable systems is generally conducted in combination with a stabilizing feedback, so our approach assumes that any necessary stabilization has been performed. The corresponding Lyapunov functions are learned by FNNs using an ISS approach. To the best of the authors’ knowledge, this is the first work analyzing the stability of errors using ISS and finding the ISS Lyapunov function with a neural network for state estimation.
We use a hybrid approach that combines both model and data. We employ a shallow JRN, presented in Section II and Section III-A, to estimate the states of the dynamical system. Data for training the recurrent network is model-based and generated as explained in Section II-B. This ensures that the training data reflect the assumptions of a Kalman filter and an extended Kalman filter. Input-to-state stability analysis of the Jordan recurrent network-based estimator is discussed in Section II, while the method to find the corresponding Lyapunov functions is explained in Section II-B. The approach is then tested in Section IV on three different discrete-time systems, obtained from the discretization of ODEs, one linear and two nonlinear. This is followed by a discussion of the convergence and input-to-state stability of the results.
II Input-to-state stability analysis
@@
where $x^{(t)}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state vector at time-step $t$, $w^{(t+1)}\in\mathbb{R}^{n}$ is the process noise vector at time-step $t+1$, $y^{(t+1)}\in\mathcal{Y}\subseteq\mathbb{R}^{m}$ is the measurement vector at time-step $t+1$, $v^{(t+1)}\in\mathbb{R}^{m}$ is the measurement noise vector at time-step $t+1$, $x^{0}\in\mathbb{R}^{n}$ is the true initial condition and $\bar{x}^{0}\in\mathbb{R}^{n}$ is the initial condition noise. Here, $\mathcal{X}$ and $\mathcal{Y}$ are the state space and measurement space, respectively, and $n\in\mathbb{N}$ is the total number of states and $m\in\mathbb{N}$ is the number of states whose noisy measurements are available. The functions $f$ and $h$ are assumed to be sufficiently smooth.
Our aim is to estimate the state vector $x^{(t+1)}$ by $\hat{x}^{(t+1)}$ based on the measurement vector $y^{(t+1)},$ previous state vector estimate $\hat{x}^{(t)}$ and the unknown noisy initial condition $x^{(0)}$.
@@
where $\sigma$ is the activation function and $\hat{x}^{(t+1)}$ is the estimated state vector at time-step $t+1$. The hidden layer vector at time-step $t+1$ is given by $a^{(t+1)}.$ The weights and biases of the network are represented by $W_{ay},W_{ax}$ and $W_{xa}$. The network is depicted in Fig. 1. This is different from an ERN because the recurrent connections for an ERN are from the previous hidden layer $a^{(t)}$ to the next hidden layer $a^{(t+1)}$, whereas for a JRN, they are from the previous output $\hat{x}^{(t)}$ to next hidden layer $a^{(t+1)}.$ Thus, for an ERN, the term $W_{ax}\hat{x}^{(t)}$ would be replaced by $W_{aa}{a}^{(t)}$ where $W_{aa}$ is a weight matrix. Note that we are considering a bias-free network.
To analyze the stability of the neural estimator, we first consider the case where there are no process noise and measurement noise in (1), i.e., $w^{(t+1)}=v^{(t+1)}=0$ for all $t$, and $\bar{x}^{0}=0$, as follows,
@@
Definition II.2
A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{KL}$-function if, for each fixed $t$, the function $\beta(\cdot,t)$ is a $\mathcal{K}$-function with respect to $r$, and for each fixed $r$, the function $\beta(r,\cdot)$ is decreasing with respect to $t$, and $\beta(r,t)\rightarrow 0$ as $t\rightarrow\infty.$
Definition II.3
The origin of (3) is globally asymptotically stable if, for each $\varepsilon>0$, there is a $\delta=\delta(\varepsilon)>0$ such that
@@
Definition II.4
[24] The discrete-time system (5) is input-to-state stable (ISS) if there exists a $\mathcal{KL}$-function $\beta$ and a $\mathcal{K}$-function $\gamma$, such that for each $t\in\mathbf{Z}_{+}$, the following inequality holds,
@@
$\begin{split}V(g(e^{(t)},x^{(t)}))-V(e^{(t)})&\leq-\alpha_{3}(\|e^{(t)}\|)+\gamma(\|x^{(t)}\|)\\ &\forall e^{(t)}\in\mathcal{E},\forall x^{(t)}\in\mathcal{X}.\end{split}$ (10)
Theorem II.6
@@
$\|x^{(t)}(\xi)\|\leq\beta_{1}(\|\xi\|,t),$ (11)
$\|e^{(t)}(\eta,x)\|\leq\beta_{2}(\|\eta\|,t)+\gamma(\|x^{(t)}\|),$ (12)
@@
Let $s^{(t)}$ denote the concatenation of the state $x^{(t)}$ and the error state $e^{(t)}$, and $\zeta$ denote the origin of this cascade system. We have $\|x^{(t)}\|\leq\|s^{(t)}\|$ and $\|e^{(t)}\|\leq\|s^{(t)}\|$, and $\|s^{(t)}\|\leq\|x^{(t)}\|+\|e^{(t)}\|$. Defining $\beta(\cdot,\cdot)=\beta_{1}(\cdot,\cdot)+\beta_{2}(\cdot,\cdot)+\gamma(\beta_{1}(\cdot,\cdot))$, this yields
@@
It can be easily verified that $\beta$ is a $\mathcal{KL}$-function, which completes the proof. ∎
@@
II-A Linear systems
Consider a linear discrete-time system defined by, for matrices $A$ and $H$,
@@ II-A Linear systems
Here, $Q$ is a symmetric positive definite matrix. In this case, it is easy to show that both properties in Def. II.5 are satisfied with $\alpha_{1}(x)=\lambda_{min}(P)x^{2}$, $\alpha_{2}(x)=\lambda_{max}(P)x^{2}$, $\alpha_{3}(x)=\frac{1}{2}\lambda_{min}(Q)x^{2}$, and $\gamma(x)=(\frac{2|\mathcal{A}^{T}P\mathcal{B}|^{2}}{\lambda_{min}(Q)}+|\mathcal{B}^{T}P\mathcal{B}|^{2})x^{2}$  [24].
II-B Nonlinear systems
For nonlinear systems, a counterexample-guided method with verification provided by SMT solvers was used to synthesize the ISS-Lyapunov functions for the error system (5).
@@ II-B Nonlinear systems
We use the square function as the activation function $\sigma$ in this network, which results in sum-of-squares (SOS)-like quadratic Lyapunov functions. Consequently, we have $V(0)=0$. As a result, a valid ISS Lyapunov function is attained when properties (9) and  (10) are satisfied. Its falsification constraints can then be written as a first-order logic formula over the real numbers. This yields
@@
where, $r_{e}$ and $r_{x}$ are the radii of the so-called valid region for the error state and the state respectively, on which we verify the condition using SMT solvers. When the SMT solver returns UNSAT, a valid ISS Lyapunov function is obtained. Otherwise, it produces counterexamples that can be added to the training dataset of the neural network for finding Lyapunov function candidates. It is worth mentioning that due to the nature of the SMT solvers, we typically need to exclude a small region around the origin when verifying the falsification constraints. Around the origin, the linearization of the nonlinear system dominates, and we can implement the method for the linearized model to provide stability guarantees as detailed in Section II-A. We refer the readers to [26] for a detailed rigorous proof and algorithm.
@@
$\begin{split}L(\theta)&=\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\max\bigg{(}0,V_{\theta}(g(e_{i},x_{j}))-V_{\theta}(e_{i})\\ &+\alpha_{3}(|e_{i}|)-\gamma(|x_{j}|)\bigg{)}+\max\bigg{(}0,V_{\theta}(e_{i})-\alpha_{2}(|e_{i}|)\bigg{)}\\ &+\max\bigg{(}0,\alpha_{1}(|e_{i}|)-V_{\theta}(e_{i})\bigg{)}.\end{split}$ (22)
This is known as the positive penalty for the violation of conditions in Definition II.5.
III Implementation
The initial conditions are sampled over an interval. The initial condition is considered to be Gaussian. The covariance of the initial condition is $(0.01)\times I$ where $I$ is an identity matrix of appropriate dimensions. Similarly, process and measurement noises are assumed to be Gaussian with zero mean and covariance $0.01\times I$ where matrix $I$ is of appropriate dimensions. This is done for each sequence in the data set. Since the approach is state estimation, knowledge of the functions $f$ and $h$ in (1) is used in data generation.
The data set is divided into three parts: training, validation and testing in the ratio $80:10:10$
Figure 2: The graph represents a comparison between JRN and KF Error$(t),0\leq t\leq 100$ values of Jordan recurrent network based state estimation and the Kalman filter based state estimation for mass-spring damper system. The magnified graph within the graph provides a closer look at the error values from time $19$ to $20$ seconds. The x-axis of both the original and the magnified graphs represents time in seconds while the y-axis represents error at time $t.$ The blue dotted line graph represents the error values for JRN based estimates while the orange solid line represents the error values for the Kalman filter based estimates.
III-A Training and validation
A custom network with forward propagation as in (2) is implemented. The weight matrix $W_{ay}$ is initialized using Xavier uniform distribution while weight matrices $W_{ya}$ and $W_{xa}$ are initialized to be (semi) orthogonal matrices using Pytorch’s $torch.nn.init$ module. The function $\sigma(z)$ is chosen to be equal to $z$ for linear systems and is chosen to be $\tanh(z)$ for nonlinear systems. The backward propagation is implemented using PyTorch’s $backward()$ function. The loss function at time-step $t+1$ for each sequence is

### v2

@@ Abstract
State estimation refers to determining the states of a dynamical system that starts from a noisy initial condition and evolves under process noise, based on noisy measurements and a known system model. For linear dynamical systems with white Gaussian noises of known mean and variance, Kalman filtering is a well-known method that leads to stable error dynamics for detectable systems. There are some non-optimal extensions to nonlinear systems. Recent work has used neural networks to develop estimators for nonlinear systems that optimize a criterion. Stability of the error dynamics is even more important than optimality. Jordan recurrent neural networks (JRNs) have a structure that mimics that of a dynamical system and are thus appealing for estimator design. We show that a JRN performs better than an extended Kalman filter(EKF) and unscented Kalman filter(UKF) for several examples. The main contribution of this paper is an input-to-state stability analysis of the error dynamics of JRNs. The stability of the error dynamics of several examples is shown.
@@ I INTRODUCTION
State estimation refers to estimating the states of a dynamical system given a model and some measurements; both of which are subject to errors and disturbances. It differs from system identification where there is no mathematical model. This is needed in some situations. However, often a model for the dynamics, although not perfect is available, and should be used in estimator design. The need for an estimator to run in real-time despite complex dynamics, inaccurate measurements, incomplete information about the initial state, and errors in the model all mean that estimation, particularly of nonlinear systems, is challenging.
Several approaches to state estimation exist dating from the early 1960s. Kalman filtering (KF) is a well known and widely used method for linear systems. If the disturbances on the model and the measurements are independent Gaussian noises, with known means and covariances then the Kalman filter yields an estimate of minimum error covariance. For nonlinear systems, several extensions of the Kalman filter are commonly used. In an extended Kalman filter (EKF), the system is linearised at each time-step and a Kalman filter for the linearized system is used [1, e.g.]. It is widely used due to its simplicity and low computational overhead. Another extension is the unscented Kalman filter (UKF). This method relies on choosing sigma points to propagate state mean and covariance [1, e.g.]. The UKF typically copes better with highly nonlinear systems, but is computationally expensive for high-order systems.
Machine learning approaches are increasingly common. A deep long short-term memory network has been used to approximate the predicted and filtered state using observations in [2]. As for optimal control, a Hamilton-Jacobi-Bellman partial differential equation needs to be solved to find the minimum variance (or maximum likelihood) estimator. Its solution is very difficult for nonlinear systems since the number of independent variables equals the order of the system (the curse of dimensionality). In [3], a neural network is used to find solutions to the Hamilton-Jacobi-Bellman equation, assuming that a smooth solution exists. In [4], a non-optimal design for discrete-time nonlinear systems is implemented using unsupervised learning. In another work, [5] compared the method of least squares, Elman recurrent neural network (ERN) and feedforward neural network (FNN) for noisy time series data with fully observable states and concluded that ERN provides the best estimation. Other papers that explore RNNs for state estimation include [6, 7, 8, 9]. Recurrent neural networks have been shown to work well for system identification[10]. Jordan recurrent networks (JRNs) have been used for system identification by several researchers and have been shown to work as well as ERNs [11, 12, 13]. In [14] the structure of a Jordan recurrent network was extended to an LSTM and compared to an Elman long short-term memory network for estimator design. The training time considerably decreased when using JLSTM for state estimation. This is consistent with the fact that the structure of a Jordan RNN is closer to that of a dynamical system than that of a Elman RNN [15].
@@
Just as stability is critical for a control system, in the absence of disturbances, the estimated state must converge to the true state. Small disturbances should increase the error by a small amount. In other words, the error dynamics must be input-to-state stable (ISS) [16, 17]. It is known that for a linear system, if the measurements satisfy the weak assumption of detectability, the error dynamics are stable. It was recently shown [18] that if a nonlinear system is locally detectable, then the error dynamics of an EKF are locally ISS. If the system is uniformly observable and can be put globally into the normal form, along with some other technical assumptions on the system dynamics then the EKF error dynamics are stable [19, sec. 2.4]. For a discrete-time UKF satisfying certain assumptions, the estimation error remains bounded [20]. The stability of errors of RNNs as state estimators has not been well explored. However, there are several approaches to establishing the stability of RNNs by finding a Lyapunov function [21, 22, 23]. This suggested extending this work to show stability of error dynamics. The issue with ERNs in this context is that proving the stability of the estimator is difficult due to its dependence on the hidden neuron state at the previous time-step. This is not the case for a JRN. We therefore design an estimator using JRNs and utilize an input-to-state stability (ISS) approach to analyze the stability of the error dynamics. Finding Lyapunov functions is generally challenging. Several papers [24, 23] have shown that Lyapunov functions can be attained and represented by expressions of the compositional structure of the neural network, the correctness being guaranteed by satisfiability modulo theories (SMT) solvers. We use this approach to establish a condition under which the error dynamics are stable if the original system is stable. Estimation of unstable systems is generally conducted in combination with a stabilizing feedback; our approach assumes that any necessary stabilization has been performed.
The use of Jordan recurrent neural network for state estimation is described in Section II. Then in Section III an ISS approach is used to provide a characterization of a Lyapunov function for the error dynamics. Data generation and training are described in Section IV. The approach is illustrated in Section V with three different discrete-time systems.
Figure 1: Jordan recurrent network (JRN) for state estimation showing output-to-hidden recurrent connections.
II Jordan recurrent neural network for state estimation
@@
where $x^{(t)}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state vector at time-step $t$, $w^{(t+1)}\in\mathbb{R}^{n}$ is the process noise vector at time-step $t+1$, $y^{(t+1)}\in\mathcal{Y}\subseteq\mathbb{R}^{m}$ is the measurement vector at time-step $t+1$, $v^{(t+1)}\in\mathbb{R}^{m}$ is the measurement noise vector at time-step $t+1$, $x^{0}\in\mathbb{R}^{n}$ is the true initial condition and $\bar{x}^{0}\in\mathbb{R}^{n}$ is the initial condition noise. Here, $\mathcal{X}$ and $\mathcal{Y}$ are the state space and measurement space, respectively, and $n\in\mathbb{N}$ is the total number of states and $m\in\mathbb{N}$ is the number of noisy measurements. The functions $f$ and $h$ are assumed to be sufficiently smooth.
Our aim is to estimate the state vector $x^{(t+1)}$ by $\hat{x}^{(t+1)}$ based on the measurement vector $y^{(t+1)}$ and the previous state vector estimate $\hat{x}^{(t)}.$
@@
where $\sigma$ is the activation function and $\hat{x}^{(t+1)}$ is the estimated state vector at time-step $t+1$. The hidden layer vector at time-step $t+1$ is given by $a^{(t+1)}.$ The weights and biases of the network are represented by $W_{ay},W_{ax}$ and $W_{xa}$. The network is depicted in Fig. 1.
This is different from an ERN because the recurrent connections for an ERN are from the previous hidden layer $a^{(t)}$ to the next hidden layer $a^{(t+1)}$, whereas for a JRN, they are from the previous output $\hat{x}^{(t)}$ to next hidden layer $a^{(t+1)}.$ Thus, for an ERN, the term $W_{ax}\hat{x}^{(t)}$ would be replaced by $W_{aa}{a}^{(t)}$ where $W_{aa}$ is a weight matrix. Note that we are considering a bias-free network.
III Input-to-state stability analysis
To analyze the stability of the neural estimator, we first consider the case with no process and measurement noise in (1), i.e., $w^{(t+1)}=v^{(t+1)}=0$ for all $t$, and $\bar{x}^{0}=0$:
@@ III Input-to-state stability analysis
Definition III.2
A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{KL}$-function if, for each $t$, $\beta(\cdot,t)$ is a $\mathcal{K}$-function with respect to $r$, and for each $r$, $\beta(r,\cdot)$ is decreasing with respect to $t$, and $\beta(r,t)\rightarrow 0$ as $t\rightarrow\infty.$
Definition III.3
The origin is globally asymptotically stable for (3) if, for each $\varepsilon>0$, there is a $\delta=\delta(\varepsilon)>0$ such that
@@
Definition III.4
[25] The discrete-time system (5) is input-to-state stable (ISS) if there exists a $\mathcal{KL}$-function $\beta$ and a $\mathcal{K}$-function $\gamma$, such that for each $t\in\mathbf{Z}_{+}$,
@@
$V(g(e^{(t)},x^{(t)}))-V(e^{(t)})$ $\leq-\alpha_{3}(\|e^{(t)}\|)+\gamma(\|x^{(t)}\|)$
$\forall e^{(t)}\in\mathcal{E},\forall x^{(t)}\in\mathcal{X}.$ (10)
Theorem III.6
@@
$\|x^{(t)}(\xi)\|$ $\leq\beta_{1}(\|\xi\|,t),$
$\|e^{(t)}(\eta,x)\|$ $\leq\beta_{2}(\|\eta\|,t)+\gamma(\|x^{(t)}\|),$
@@
Let $s^{(t)}$ denote the concatenation of the state $x^{(t)}$ and the error state $e^{(t)}$, and $\zeta$ denote the origin of this cascade system. We have $\|x^{(t)}\|\leq\|s^{(t)}\|$ and $\|e^{(t)}\|\leq\|s^{(t)}\|$, and $\|s^{(t)}\|\leq\|x^{(t)}\|+\|e^{(t)}\|$ (because for any 2 non-negative real numbers $a$ and $b,\sqrt{a^{2}+b^{2}}\leq a+b$). Defining $\beta(\cdot,\cdot)=\beta_{1}(\cdot,\cdot)+\beta_{2}(\cdot,\cdot)+\gamma(\beta_{1}(\cdot,\cdot))$, this yields
@@
It can be easily verified that $\beta$ is a $\mathcal{KL}$-function. ∎
@@
Linear Systems. Consider a linear discrete-time system defined by, for matrices $A$ and $H$,
@@
Here, $Q$ is a symmetric positive definite matrix. In this case, it is easy to show that both properties in Def. III.5 are satisfied with $\alpha_{1}(x)=\lambda_{min}(P)x^{2}$, $\alpha_{2}(x)=\lambda_{max}(P)x^{2}$, $\alpha_{3}(x)=\frac{1}{2}\lambda_{min}(Q)x^{2}$, and $\gamma(x)=(\frac{2|\mathcal{A}^{T}P\mathcal{B}|^{2}}{\lambda_{min}(Q)}+|\mathcal{B}^{T}P\mathcal{B}|^{2})x^{2}$  [25].
Nonlinear systems. For nonlinear systems, a counterexample-guided method with verification provided by SMT solvers was used to synthesize the ISS-Lyapunov functions for the error system (5).
@@
We use the square function as the activation function $\sigma$ in this network, which results in sum-of-squares (SOS)-like quadratic Lyapunov functions. Then $V(0)=0$. As a result, a valid ISS Lyapunov function is attained when properties (9) and  (10) are satisfied. Its falsification constraints can then be written as a first-order logic formula over the real numbers. This yields
@@
where, $r_{e}$ and $r_{x}$ are the radii of the so-called valid region for the error state and the state respectively, on which we verify the condition using SMT solvers. When the SMT solver returns UNSAT, a valid ISS Lyapunov function is obtained. Otherwise, it produces counterexamples that can be added to the training dataset of the neural network for finding Lyapunov function candidates. Due to the nature of the SMT solvers, we typically need to exclude a small region around the origin when verifying the falsification constraints. Around the origin, the linearization of the nonlinear system dominates, and we implement the method for the linearized model, described above, to provide stability guarantees. We refer the readers to [21] for a detailed proof and algorithm.
@@
$L(\theta)$ $=\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\max\bigg{(}0,V_{\theta}(g(e_{i},x_{j}))-V_{\theta}(e_{i})$
$+\alpha_{3}(|e_{i}|)-\gamma(|x_{j}|)\bigg{)}+\max\bigg{(}0,V_{\theta}(e_{i})-\alpha_{2}(|e_{i}|)\bigg{)}$
$+\max\bigg{(}0,\alpha_{1}(|e_{i}|)-V_{\theta}(e_{i})\bigg{)}.$ (20)
This is known as the positive penalty for the violation of conditions in Definition III.5.
IV Implementation
The datasets are obtained by discretizing common ordinary differential equations. We use a zero order hold discretization for linear systems. For nonlinear continuous-time state space systems, we use the RK-45 discretization using Python’s $scipy.integrate.solve\_ivp$ function. The initial condition is considered to be Gaussian with mean sampled uniformly from the interval $[-1,1]\times[-1,1]\subset\mathbb{R}\times\mathbb{R}$ for each sequence and known covariance. Process and measurement noises are assumed to be Gaussian with zero mean and known covariance. For the systems under consideration, the above 3 covariance matrices are assumed to be $0.01\times I$ where $I$ is an identity matrix of appropriate dimensions. Also, knowledge of the functions $f$ and $h$ in (1) is used to generate data.
For the linear system, we consider a total of $100$ sequences and $200$ for non-linear systems. The data set is divided into three parts: training, validation and testing in the ratio $80:10:10$ sequences. Each of the generated sequences is independent of the other sequences in the dataset.
A custom network with forward propagation as in (2) is implemented. The weight matrix $W_{ay}$ is initialized using Xavier uniform distribution while weight matrices $W_{ya}$ and $W_{xa}$ are initialized to be (semi) orthogonal matrices using Pytorch’s $torch.nn.init$ module. The function $\sigma(z)$ is chosen to be equal to $z$ for the linear system and $\tanh(z)$ for the nonlinear systems. The backward propagation is implemented using PyTorch’s $backward()$ function. We consider the mean squared error loss function at time-step $t+1$ for each sequence